# Channel-Imposed Fusion: A Simple yet Effective Method for Medical Time Series Classification

### v1

@@ Appendix G Limitations and Future Work
Furthermore, integrating domain-specific prior knowledge into medical time-series analysis can reveal and model relationships between channels more precisely. By combining medical expertise, clinical experience, and existing pathological data, we can enhance model interpretability and predictive performance, supporting clinical decision-making and interventions. Future research could focus on incorporating such prior knowledge to advance intelligent medical analytics.
Finally, we must acknowledge that the development trends in the field of artificial intelligence highlight the importance of architectural innovation. Future research should focus on designing novel architectures that align more closely with the CIF method, combining the strengths of existing models. For example, the local feature extraction capabilities of CNNs LeCun et al. [1989], the temporal stability of TCNs Bai et al. [2018] for long sequences, the long-term dependency modeling of RNNs Rumelhart et al. [1986] and LSTMs Hochreiter and Schmidhuber [1997], the global modeling efficiency of Transformers Vaswani et al. [2017], the resource-efficient computation of Mamba Gu and Dao [2023], and the hybrid recurrence-attention structure of RWKV Peng et al. [2023]. By adapting and integrating these methods, we aim to build a powerful model that is not only deeply compatible with the CIF framework but also capable of efficiently handling complex medical time-series data.
Appendix H Broader Impacts
In this study, we take the first step toward a paradigm shift in medical time-series analysis, transitioning from a focus on model architecture improvements to emphasizing the intrinsic structure of the data. Our experimental results, which outperform several state-of-the-art methods across multiple datasets, validate the effectiveness of this approach. We emphasize that uncovering the essential characteristics in medical time-series data is a valuable research direction, particularly given the critical need for model interpretability in medical applications. However, practitioners should be cognizant of the method’s limitations.

### v2

@@ Appendix G Limitations and Future Work
Furthermore, incorporating domain-specific prior knowledge into medical time-series analysis can more precisely reveal and model relationships between channels. By integrating medical expertise, clinical experience, and existing pathological data, the interpretability and predictive performance of models can be enhanced, thereby supporting clinical decision-making and interventions. On this basis, frequency-domain analysis Hu et al. (2025); Nason & Sachs (1999); Yi et al. (2025) offers an additional perspective: by applying Fourier transform or wavelet decomposition to the signals, physiological features at different frequency components can be identified, revealing patterns that are difficult to capture in the time domain. This is particularly valuable for noise reduction, extraction of periodic signals, and detection of pathological events, and can also provide richer feature representations for model inputs. Future research could further explore how to combine time-domain and frequency-domain information, integrating domain priors to improve the accuracy and robustness of intelligent medical analytics.
Finally, we must acknowledge that the development trends in the field of artificial intelligence highlight the importance of architectural innovation. Future research should focus on designing novel architectures that align more closely with the CIF method, combining the strengths of existing models. For example, the local feature extraction capabilities of CNNs LeCun et al. (1989), the temporal stability of TCNs Bai et al. (2018) for long sequences, the long-term dependency modeling of RNNs Rumelhart et al. (1986) and LSTMs Hochreiter & Schmidhuber (1997), the global modeling efficiency of Transformers Vaswani et al. (2017), the resource-efficient computation of Mamba Gu & Dao (2023), and the hybrid recurrence-attention structure of RWKV Peng et al. (2023). By adapting and integrating these methods, we aim to build a powerful model that is not only deeply compatible with the CIF framework but also capable of efficiently handling complex medical time-series data.